a child of mass 25 kg sits at rest at the top of a frictionless slide 2.0m off the ground. He then slides down. what is his kinetic energy at the bottom of the slide?
Answer #1

v = \sqrt{}2gh = \sqrt{}2x9.8x2.0 = 6.26m/s.

kinetic energy = mv2 / 2 = 25x(6.26)2 / 2 = 489.84J.